OpenStudy (anonymous):

What is the value of x? x/6= 15/18 A.3 B.5 C.45 D.90

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

5*3=15 6*3=18 Thus 15/18 x=5

OpenStudy (cggurumanjunath):